# Exact Mobility Edges in a Disorder-Free Dimerized Stark Lattice with Effective Unbounded Hopping

### v2

@@ Acknowledgements.
The authors would like to thank Peijie Chang and Dan Long for their helpful discussions. This work is supported by the National Natural Science Foundation of China under Grant No. 62131002.
@@ Appendix A Gauge Transformation to Real Symmetric Form
Recall that the effective model derived in the main text possesses a real inter-cell hopping $J_{\text{inter}}=t_{2}$ and a complex intra-cell hopping
@@
We define a site-dependent gauge transformation via the unitary operator $S$, which acts on the basis states as $|A_{n}\rangle\to e^{i\Phi_{n,A}}|A_{n}\rangle$ and $|B_{n}\rangle\to e^{i\Phi_{n,B}}|B_{n}\rangle$. To render the intra-cell hopping real, the phase difference within a cell must satisfy $\Phi_{n,B}-\Phi_{n,A}=\theta_{n}$. Simultaneously, to keep the inter-cell hopping $t_{2}$ (connecting $B_{n}$ and $A_{n+1}$) real, we require $\Phi_{n+1,A}=\Phi_{n,B}$.
@@
$S=\text{diag}(e^{-i\Phi_{-N,A}},e^{-i\Phi_{-N,B}},\cdots,e^{-i\Phi_{N,A}},e^{-i\Phi_{N,B}})$ (19)
@@ Appendix B Restrictions on the Spectrum of Jacobi Matrices
We first cite Theorem 2.1 in the original work of Simon and Spencer [SimonSpencer_1989], which forbids the absolutely continuous (AC) spectrum with unbounded potentials. In Theorem 2.1, they assume the operator $h$ to be
@@
leads to empty AC spectrum. From Eq. (21) we clearly see the hopping term is assumed to be nearest-neighbor (NN) and uniform ($u(n\pm 1)$ has coefficient 1). Our effective model violates the uniformity condition ($a_{n}\sim n$). Furthermore, our original model Eq. (1) violates the NN condition directly due to the sublattice spinor structure. Thus, the $n$-dependent hopping helps evade the restriction from the Simon-Spencer theorem.
Next, we examine broader conditions for the existence or absence of spectral components in unbounded Jacobi matrices in mathematical physics. We will demonstrate that our effective Hamiltonian $H^{\prime\prime}$ does not satisfy the sufficient conditions for an empty AC spectrum or empty pure point (PP) spectrum found in the relevant literature [Cojuhari_2008, Swiderski_2016].
@@
$a_{2n}$ $=t_{2},\quad$ $a_{2n+1}$ $=\sqrt{t_{1}^{2}+F^{2}n^{2}/4},$ (23)
$b_{2n}$ $=-\frac{Fn}{2},\quad$ $b_{2n+1}$ $=-\frac{Fn}{2}.$
Theorem 2.2(c) of Ref. [Cojuhari_2008] provides a sufficient condition for the absence of an AC spectrum. The spectrum is empty if
$\liminf_{n\to\infty}a_{n}>0,$ $\quad\sup_{n}\frac{b_{n}^{2}+b_{n+1}^{2}}{a_{n}^{2}}<\infty,$ (24)
$\liminf_{n\to\infty}$ $(1-\frac{b_{n}b_{n+1}}{a_{n}^{2}})>0.$
@@
Regarding the PP spectrum, Ref. [Swiderski_2016] [Theorem 1.1] provides conditions forbidding the PP spectrum of a Jacobi matrix. A necessary condition for this theorem to hold is
@@
In summary, the staggered nature of the weights $\{a_{n}\}$—alternating between bounded ($t_{2}$) and unbounded ($J_{\text{intra}}\sim n$)—is the critical feature. It places the model in a “gap” between theorems that require uniform or bounded hopping (Simon-Spencer) and theorems designed for uniformly growing matrices. We direct readers interested in the mathematical construction of other operators supporting mobility edges (MEs) to Ref. [Janas_2003] and Ref. [Janas_2009].
@@ Appendix C Derivation of the Spectrum
$I(k)$ $=\int^{k}\frac{t_{1}^{2}+t_{2}^{2}+2t_{1}t_{2}\cos q-E^{2}}{E-t_{2}\sin q}dq$ (27)
$=(t_{1}^{2}+t_{2}^{2}-E^{2})\int^{k}\frac{dq}{E-t_{2}\sin q}+\int^{k}\frac{2t_{1}t_{2}\cos q}{E-t_{2}\sin q}dq.$
@@
$\int\frac{dq}{\alpha-\sin q}$ $=\int\frac{2dx}{\alpha(1+x^{2})-2x}$ (28)
$=\frac{1}{\sqrt{1-\alpha^{2}}}\ln\left(\frac{\alpha\tan(k/2)-1-\sqrt{1-\alpha^{2}}}{\alpha\tan(k/2)-1+\sqrt{1-\alpha^{2}}}\right).$
@@
This corresponds to Eq. (6) in the main text. To ensure the wavefunction is single-valued in the Brillouin Zone (BZ), we impose the condition $\phi_{B}(\pi)=\phi_{B}(-\pi)$. While the function is periodic, the logarithmic term in Eq. (29) accumulates a phase of $2\pi$ when traversing the BZ. Specifically, the argument of the logarithm has two branches; the branch with the positive imaginary part spans the range $\pi$ to $\pi$, while the negative branch spans the range $2\pi$ to $\pi$. This $2\pi$ phase winding enforces the exponent to be an integer. Note that for localized states $|E|>t_{2}$, the term $\sqrt{1-\alpha^{2}}$ becomes purely imaginary, rendering the prefactor of the logarithm real
@@
Since $(m^{2}F^{2}/2+t_{1}^{2})^{2}=m^{4}F^{4}/4+m^{2}t_{1}^{2}F^{2}+t_{1}^{4}$, the term in the parentheses is negative. Since $m<0$, the prefactor is negative, making the total derivative positive. Thus, $E_{m}^{2}$ increases with $m$ (decreases as $m\to$-$\infty$). The asymptotic limit is
$E_{m}^{2}\xrightarrow{m\to-\infty}t_{2}^{2}+\frac{t_{1}^{4}}{m^{2}F^{2}}.$ (36)
@@ Appendix D Large-Scale Numerical Methods
    Shift-Invert Spectral Transformation: To target interior eigenvalues (e.g., the continuum states near $E=0.5t_{2}$), we solve the transformed eigenvalue problem $(H-\sigma I)^{-1}\psi=\nu\psi$, where $\sigma$ is the target energy shift [Saad_2011]. The state closest to $\sigma$ corresponds to the eigenvalue $\nu$ with the largest magnitude, which converges rapidly in the Krylov iteration. To track the evolution of specific states, we set $\sigma=E_{m}^{\text{theo}}$ for discrete localized levels (fixed index $m$) and $\sigma=0.5t_{2}$ for the extended continuum, ensuring we probe bulk properties away from edge effects.
@@
    Banded Linear Solver: Under open boundary conditions (OBCs), the effective Hamiltonian $H^{\prime\prime}$ (derived via the gauge transformation) is tridiagonal. Instead of general sparse formats, we store $H$ in a packed banded format. The linear inversion step $(H-\sigma I)x=b$ is performed using the Thomas algorithm (via LAPACK’s gtsv routine), which requires only $O(L)$ time and memory, avoiding the fill-in associated with general sparse LU decomposition.
@@ E.1 Accuracy Comparison
We validate our disk-backed solver against the standard scipy.sparse.linalg.eigsh implementation for system sizes where the latter fits in memory ($L\approx 10^{3}$ to $10^{6}$). Fig. E1(a) shows the relative error of energy $|\Delta E|/|E|$ for the three representative states used in the main text. We repeat the calculation 10 times and report the maximum relative error in this figure. For the discrete levels, the error is close to machine precision for both branches while it is more significant (with maximum value $\sim 10^{-13}$) for the continuum band. The larger relative error observed in the continuum branch is a signature of the extremely dense spectrum, where the eigenvalue conditioning deteriorates rapidly with size and results in numerical instability of the solution. Note that an offset of $10^{-20}$ was added to the plot to handle exact zero differences on the logarithmic scale.
@@ E.3 Statistical Universality
To verify that the single state tracked in the main text ($E\approx 0.5t_{2}$) is representative of the entire extended phase, we calculate the inverse participation ratio (IPR) for 20 randomly sampled states within the extended continuum for each system size. Fig. E2 displays the mean IPR (red lines) and the standard deviation (shaded region). The variance is minimal, and the IPR of the representative state lies within the $\pm\sigma$ band. This suggests that the $L^{-1}$ scaling is a universal bulk property of the continuum, robust against microscopic energy variations.
@@ F.1 Rotating Frame Transformation
Consider a ring resonator supporting a set of frequency modes. In the absence of group velocity dispersion (GVD), the eigenfrequencies are equidistant: $\omega_{n}=\omega_{0}+n\Omega_{\text{FSR}}$, where $n$ is the mode index relative to the carrier frequency $\omega_{0}$, and $\Omega_{\text{FSR}}$ is the free spectral range (FSR). The bare Hamiltonian in the laboratory frame is
$H_{0}=\sum_{n}(\omega_{A}+n\Omega_{A})a_{n}^{\dagger}a_{n}+(\omega_{B}+n\Omega_{B})b_{n}^{\dagger}b_{n},$ (37)
@@
$H_{h}(t)=2t_{h}\cos(\omega_{\text{mod}}t-\pi/2)\sum_{n,m}(a_{n+m}^{\dagger}a_{n}+a_{n}^{\dagger}a_{n+m})$ (38)
$+2t_{h}\cos(\omega_{\text{mod}}t+\pi/2)\sum_{n,m}(b_{n+m}^{\dagger}b_{n}+b^{\dagger}_{n}b_{n+m}),$
where $t_{h}$ is the hopping strength proportional to the radio frequency (RF) drive amplitude. We apply two coherent RF drives with the same hopping strength on the two resonators and a $\pi$ phase difference, yielding the $\pm\pi/2$ in the phases. Similarly, the direct current (DC) bias and RF drive on the Mach-Zehnder Interferometer (MZI) contribute to the coupling between the two resonators. The DC bias couples modes with the same frequencies (vertical coupling)
@@
$U(t)=\exp\left[-it\sum_{n}(\omega_{A}+n\omega_{\text{mod}})(a_{n}^{\dagger}a_{n}+b_{n}^{\dagger}b_{n})\right].$ (41)
@@
$i\dot{U}^{\dagger}U=-\sum_{n}\left[(\omega_{A}+n\omega_{\text{mod}})(a_{n}^{\dagger}a_{n}+b^{\dagger}_{n}b_{n})\right].$ (42)
@@
$H_{\text{diag}}$ $=\sum_{n}n\left[(\Omega_{A}-\omega_{\text{mod}})a_{n}^{\dagger}a_{n}+(\Omega_{B}-\omega_{\text{mod}})b_{n}^{\dagger}b_{n}\right]$ (43)
$+\sum_{n}(\omega_{B}-\omega_{A})b^{\dagger}_{n}b_{n}.$
This derivation explicitly shows that detuning the resonator’s FSR from the modulation frequency generates a linear Stark potential $V_{n}=\Delta n$ while the difference in the central frequencies leads to a constant shift on the B-sublattice. We define the detuning parameters $\Delta_{A(B)}=\Omega_{A(B)}-\omega_{\text{mod}}$ and $\delta_{BA}=\omega_{B}-\omega_{A}$. By designing slightly different FSRs for the two resonators, we set $\Delta_{A}\approx 0$ (resonance) and $\Delta_{B}=F\neq 0$ (detuned) to realize the sublattice-selective potential.
@@
$U^{\dagger}a_{n+m}^{\dagger}a_{n}U$ $=U_{m+n}^{\dagger}a_{n+m}^{\dagger}U_{m+n}U_{n}^{\dagger}a_{n}U_{n}$ (44)
$=e^{im\omega_{\text{mod}}t}a_{n+m}^{\dagger}a_{n},$
where $U_{n}=\exp\left[-it(\omega_{A}+n\omega_{\text{mod}})(a_{n}^{\dagger}a_{n}+b_{n}^{\dagger}b_{n})\right]$. Similarly, we have
$U^{\dagger}b_{n+m}^{\dagger}a_{n}U=e^{im\omega_{\text{mod}}t}b_{n+m}^{\dagger}a_{n},$ (45)
$U^{\dagger}a_{n+m}^{\dagger}b_{n}U=e^{im\omega_{\text{mod}}t}a_{n+m}^{\dagger}b_{n}.$
Expanding the cosine modulation as $\left[e^{i(\omega_{\text{mod}}t+\phi)}+e^{-i(\omega_{\text{mod}}t+\phi)}\right]/2$, and applying the Rotating Wave Approximation (RWA) to discard the rapidly oscillating terms, we obtain the static coupling
$U^{\dagger}H_{h}U$ $\approx it_{h}\sum_{n}(a_{n+1}^{\dagger}a_{n}+b_{n}^{\dagger}b_{n+1}-a_{n}^{\dagger}a_{n+1}-b_{n+1}^{\dagger}b_{n}),$ (46)
$U^{\dagger}H_{c}U$ $\approx t_{c}\sum_{n}(a^{\dagger}_{n+1}b_{n}+b^{\dagger}_{n}a_{n+1}+b_{n+1}^{\dagger}a_{n}+a^{\dagger}_{n}b_{n+1}),$
$U^{\dagger}H_{v}U$ $=H_{v}.$
Thus, the Hamiltonian under time-dependent modulation maps to a static lattice in the frequency dimension in the rotating frame, realizing the model described by Eq. (1) with the addition of the potential shift $\sum_{n}\delta_{BA}b^{\dagger}_{n}b_{n}$ .

### v3

@@ Acknowledgements.
The authors thank Peijie Chang and Dan Long for their helpful discussions. This work is supported by the National Natural Science Foundation of China under Grant No. 62131002.
@@ Appendix A Gauge Transformation to Real Symmetric Form
Recall that the effective model derived in the main text possesses a real intercell hopping $J_{\text{inter}}=t_{2}$ and a complex intracell hopping
@@
We define a site-dependent gauge transformation via the unitary operator $S$, which acts on the basis states as $|A_{n}\rangle\to e^{i\Phi_{n,A}}|A_{n}\rangle$ and $|B_{n}\rangle\to e^{i\Phi_{n,B}}|B_{n}\rangle$. To render the intracell hopping real, the phase difference within a cell must satisfy $\Phi_{n,B}-\Phi_{n,A}=\theta_{n}$. Simultaneously, to keep the intercell hopping $t_{2}$ (connecting $B_{n}$ and $A_{n+1}$) real, we require $\Phi_{n+1,A}=\Phi_{n,B}$.
@@
$S=\text{diag}(e^{-i\Phi_{-N,A}},e^{-i\Phi_{-N,B}},\cdots,e^{-i\Phi_{N,A}},e^{-i\Phi_{N,B}}).$ (19)
@@ Appendix B Restrictions on the Spectrum of Jacobi Matrices
We first cite Theorem 2.1 in the original work of Simon and Spencer [63], which forbids the AC spectrum with unbounded potentials. In Theorem 2.1, they assume the operator $h$ to be
@@
leads to an empty AC spectrum. From Eq. (21) we clearly see the hopping term is assumed to be NN and uniform $u(n\pm 1)$ has coefficient 1]. Our effective model violates the uniformity condition ($a_{n}\sim n$). Furthermore, our original model Eq. (1) violates the NN condition directly due to the sublattice spinor structure. Thus, the $n$-dependent hopping helps evade the restriction from the Simon-Spencer theorem.
Next, we examine broader conditions for the existence or absence of spectral components in unbounded Jacobi matrices in mathematical physics. We will demonstrate that our effective Hamiltonian $H^{\prime\prime}$ does not satisfy the sufficient conditions for an empty AC spectrum or empty pure point (PP) spectrum found in the relevant literature [52, 65].
@@
$aligned a_{2n}&=t_{2},\quad&a_{2n+1}&=\sqrt{t_{1}^{2}+F^{2}n^{2}/4},\\ b_{2n}&=-\frac{Fn}{2},\quad&b_{2n+1}&=-\frac{Fn}{2}.$ (23)
Theorem 2.2(c) of Ref. [52] provides a sufficient condition for the absence of an AC spectrum. The spectrum is empty if
[…]
@@
Regarding the PP spectrum, Ref. [65, Theorem 1.1] provides conditions forbidding the PP spectrum of a Jacobi matrix. A necessary condition for this theorem to hold is
@@
In summary, the staggered nature of the weights $\{a_{n}\}$—alternating between bounded ($t_{2}$) and unbounded ($J_{\text{intra}}\sim n$)—is the critical feature. It places the model in a “gap” between theorems that require uniform or bounded hopping (Simon-Spencer) and theorems designed for uniformly growing matrices. We direct readers interested in the mathematical construction of other operators supporting MEs to Refs. [3435].
@@ Appendix C Derivation of the Spectrum
$aligned I(k)&=\int^{k}\frac{t_{1}^{2}+t_{2}^{2}+2t_{1}t_{2}\cos q-E^{2}}{E-t_{2}\sin q}dq\\ &=(t_{1}^{2}+t_{2}^{2}-E^{2})\int^{k}\frac{dq}{E-t_{2}\sin q}+\int^{k}\frac{2t_{1}t_{2}\cos q}{E-t_{2}\sin q}dq.$ (27)
@@
$aligned\int\frac{dq}{\alpha-\sin q}&=\int\frac{2dx}{\alpha(1+x^{2})-2x}\\ &=\frac{1}{\sqrt{1-\alpha^{2}}}\ln\left(\frac{\alpha\tan(k/2)-1-\sqrt{1-\alpha^{2}}}{\alpha\tan(k/2)-1+\sqrt{1-\alpha^{2}}}\right).$ (28)
@@
This corresponds to Eq. (6) in the main text. To ensure the wavefunction is single valued in the BZ, we impose the condition $\phi_{B}(\pi)=\phi_{B}(-\pi)$. While the function is periodic, the logarithmic term in Eq. (29) accumulates a phase of $2\pi$ when traversing the BZ. Specifically, the argument of the logarithm has two branches; the branch with the positive imaginary part spans the range $\pi$ to $\pi$, while the negative branch spans the range $2\pi$ to $\pi$. This $2\pi$ phase winding enforces the exponent to be an integer. Note that for localized states $|E|>t_{2}$, the term $\sqrt{1-\alpha^{2}}$ becomes purely imaginary, rendering the prefactor of the logarithm real
@@
Since $(m^{2}F^{2}/2+t_{1}^{2})^{2}=m^{4}F^{4}/4+m^{2}t_{1}^{2}F^{2}+t_{1}^{4}$, the term in parentheses is negative. Since $m<0$, the prefactor is negative, making the total derivative positive. Thus, $E_{m}^{2}$ increases with $m$ (decreases as $m\to$-$\infty$). The asymptotic limit is
[…]
@@ Appendix D Large-Scale Numerical Methods
    Shift-invert spectral transformation: To target interior eigenvalues (e.g., the continuum states near $E=0.5t_{2}$), we solve the transformed eigenvalue problem $(H-\sigma I)^{-1}\psi=\nu\psi$, where $\sigma$ is the target energy shift [60]. The state closest to $\sigma$ corresponds to the eigenvalue $\nu$ with the largest magnitude, which converges rapidly in the Krylov iteration. To track the evolution of specific states, we set $\sigma=E_{m}^{\text{theo}}$ for discrete localized levels (fixed index $m$) and $\sigma=0.5t_{2}$ for the extended continuum, ensuring we probe bulk properties away from edge effects.
@@
    Banded linear solver: Under OBCs, the effective Hamiltonian $H^{\prime\prime}$ (derived via the gauge transformation) is tridiagonal. Instead of general sparse formats, we store $H$ in a packed banded format. The linear inversion step $(H-\sigma I)x=b$ is performed using the Thomas algorithm (via LAPACK’s gtsv routine), which requires only $O(L)$ time and memory, avoiding the fill-in associated with general sparse $LU$ decomposition.
@@ E.1 Accuracy comparison
We validate our disk-backed solver against the standard scipy.sparse.linalg.eigsh implementation for system sizes where the latter fits in memory ($L\approx 10^{3}$ to $10^{6}$). Figure E1(a) shows the relative error of energy $|\Delta E|/|E|$ for the three representative states used in the main text. We repeat the calculation 10 times and report the maximum relative error in this figure. For the discrete levels, the error is close to machine precision for both branches while it is more significant (with maximum value $\sim 10^{-13}$) for the continuum band. The larger relative error observed in the continuum branch is a signature of the extremely dense spectrum, where the eigenvalue conditioning deteriorates rapidly with size and results in numerical instability of the solution. Note that an offset of $10^{-20}$ was added to the plot to handle exact zero differences on the logarithmic scale.
@@ E.3 Statistical universality
To verify that the single state tracked in the main text ($E\approx 0.5t_{2}$) is representative of the entire extended phase, we calculate the IPR for 20 randomly sampled states within the extended continuum for each system size. Figure E2 displays the mean IPR (red line) and the standard deviation (shaded region). The variance is minimal, and the IPR of the representative state lies within the $\pm\sigma$ band. This suggests that the $L^{-1}$ scaling is a universal bulk property of the continuum, robust against microscopic energy variations.
@@ F.1 Rotating frame transformation
Consider a ring resonator supporting a set of frequency modes. In the absence of group velocity dispersion, the eigenfrequencies are equidistant: $\omega_{n}=\omega_{0}+n\Omega_{\text{FSR}}$, where $n$ is the mode index relative to the carrier frequency $\omega_{0}$, and $\Omega_{\text{FSR}}$ is the FSR. The bare Hamiltonian in the laboratory frame is
$H_{0}=\sum_{n}(\omega_{\text{A}}+n\Omega_{\text{A}})a_{n}^{\dagger}a_{n}+(\omega_{\text{B}}+n\Omega_{\text{B}})b_{n}^{\dagger}b_{n},$ (37)
@@
$aligned H_{h}(t)=2t_{h}\cos(\omega_{\text{mod}}t-\pi/2)\sum_{n,m}(a_{n+m}^{\dagger}a_{n}+a_{n}^{\dagger}a_{n+m})\\ +2t_{h}\cos(\omega_{\text{mod}}t+\pi/2)\sum_{n,m}(b_{n+m}^{\dagger}b_{n}+b^{\dagger}_{n}b_{n+m}),$ (38)
where $t_{h}$ is the hopping strength proportional to the RF drive amplitude. We apply two coherent RF drives with the same hopping strength on the two resonators and a $\pi$ phase difference, yielding the $\pm\pi/2$ in the phases. Similarly, the DC bias and RF drive on the MZI contribute to the coupling between the two resonators. The DC bias couples modes with the same frequencies (vertical coupling)
@@
$U(t)=\exp\left[-it\sum_{n}(\omega_{\text{A}}+n\omega_{\text{mod}})(a_{n}^{\dagger}a_{n}+b_{n}^{\dagger}b_{n})\right].$ (41)
@@
$i\dot{U}^{\dagger}U=-\sum_{n}\left[(\omega_{\text{A}}+n\omega_{\text{mod}})(a_{n}^{\dagger}a_{n}+b^{\dagger}_{n}b_{n})\right].$ (42)
@@
$aligned H_{\text{diag}}&=\sum_{n}n\left[(\Omega_{\text{A}}-\omega_{\text{mod}})a_{n}^{\dagger}a_{n}+(\Omega_{\text{B}}-\omega_{\text{mod}})b_{n}^{\dagger}b_{n}\right]\\ &+\sum_{n}(\omega_{\text{B}}-\omega_{\text{A}})b^{\dagger}_{n}b_{n}.$ (43)
This derivation explicitly shows that detuning the resonator’s FSR from the modulation frequency generates a linear Stark potential $V_{n}=\Delta n$ while the difference in the central frequencies leads to a constant shift on the B-sublattice. We define the detuning parameters $\Delta_{\text{A$B$}}=\Omega_{\text{A$B$}}-\omega_{\text{mod}}$ and $\delta_{\text{BA}}=\omega_{\text{B}}-\omega_{\text{A}}$. By designing slightly different FSRs for the two resonators, we set $\Delta_{\text{A}}\approx 0$ (resonance) and $\Delta_{\text{B}}=F\neq 0$ (detuned) to realize the sublattice-selective potential.
@@
$aligned U^{\dagger}a_{n+m}^{\dagger}a_{n}U&=U_{m+n}^{\dagger}a_{n+m}^{\dagger}U_{m+n}U_{n}^{\dagger}a_{n}U_{n}\\ &=e^{im\omega_{\text{mod}}t}a_{n+m}^{\dagger}a_{n},$ (44)
where $U_{n}=\exp\left[-it(\omega_{\text{A}}+n\omega_{\text{mod}})(a_{n}^{\dagger}a_{n}+b_{n}^{\dagger}b_{n})\right]$. Similarly, we have
$aligned U^{\dagger}b_{n+m}^{\dagger}a_{n}U=e^{im\omega_{\text{mod}}t}b_{n+m}^{\dagger}a_{n},\\ U^{\dagger}a_{n+m}^{\dagger}b_{n}U=e^{im\omega_{\text{mod}}t}a_{n+m}^{\dagger}b_{n}.$ (45)
Expanding the cosine modulation as $\left[e^{i(\omega_{\text{mod}}t+\phi)}+e^{-i(\omega_{\text{mod}}t+\phi)}\right]/2$, and applying the rotating wave approximation to discard the rapidly oscillating terms, we obtain the static coupling
$aligned U^{\dagger}H_{h}U&\approx it_{h}\sum_{n}(a_{n+1}^{\dagger}a_{n}+b_{n}^{\dagger}b_{n+1}-a_{n}^{\dagger}a_{n+1}-b_{n+1}^{\dagger}b_{n}),\\ U^{\dagger}H_{c}U&\approx t_{c}\sum_{n}(a^{\dagger}_{n+1}b_{n}+b^{\dagger}_{n}a_{n+1}+b_{n+1}^{\dagger}a_{n}+a^{\dagger}_{n}b_{n+1}),\\ U^{\dagger}H_{v}U&=H_{v}.$ (46)
Thus, the Hamiltonian under time-dependent modulation maps to a static lattice in the frequency dimension in the rotating frame, realizing the model described by Eq. (1) with the addition of the potential shift $\sum_{n}\delta_{\text{BA}}b^{\dagger}_{n}b_{n}$ .